# InstructBioMol: Advancing Biomolecule Understanding and Design Following Human Instructions

### v1

@@ Abstract
Understanding and designing biomolecules, such as proteins and small molecules, is central to advancing drug discovery, synthetic biology, and enzyme engineering. Recent breakthroughs in Artificial Intelligence (AI) have revolutionized biomolecular research, achieving remarkable accuracy in biomolecular prediction and design. However, a critical gap remains between AI’s computational power and researchers’ intuition, using natural language to align molecular complexity with human intentions. Large Language Models (LLMs) have shown potential to interpret human intentions, yet their application to biomolecular research remains nascent due to challenges including specialized knowledge requirements, multimodal data integration, and semantic alignment between natural language and biomolecules. To address these limitations, we present InstructBioMol, a novel LLM designed to bridge natural language and biomolecules through a comprehensive any-to-any alignment of natural language, molecules, and proteins. This model can integrate multimodal biomolecules as input, and enable researchers to articulate design goals in natural language, providing biomolecular outputs that meet precise biological needs. Experimental results demonstrate InstructBioMol can understand and design biomolecules following human instructions. Notably, it can generate drug molecules with a 10% improvement in binding affinity and design enzymes that achieve an ESP Score of 70.4, making it the only method to surpass the enzyme-substrate interaction threshold of 60.0 recommended by the ESP developer. This highlights its potential to transform real-world biomolecular research.
@@ Introduction
Understanding and designing biomolecules is fundamental to natural science research. Biomolecules, such as proteins and small molecules, play essential roles in biological processes, and their precise manipulation is key to advancements in drug discovery, synthetic biology, and enzyme engineering [1, 2, 3]. Recent Artificial Intelligence (AI) breakthroughs have transformed research in these areas [4, 5]. Tools like AlphaFold3 [4] and RoseTTAFold All-Atom [5] have revolutionized biomolecular structure prediction, offering unprecedented accuracy and speed. Despite these advancements, a crucial challenge persists: how to effectively understand biomolecules using natural language and design them according to human intentions. This presents a gap between AI’s computational capacity and researchers’ needs to apply them to real-world problems. Consider the scenario of a researcher tasked with designing a new drug to target a protein involved in a complex disease. Traditionally, this process involves navigating vast amounts of biomolecular data, interpreting biological and chemical relationships, and iterating through trial-and-error to engineer molecules with specific properties. While AI has enhanced many aspects of this workflow, current tools often struggle to align molecular complexity with intuitive, human-driven goals articulated in natural language.
@@
Notably, Large Language Models (LLMs) [6] have demonstrated their impressive capacity to understand human intention and generate human-like responses helpfully and safely [7, 8]. This capability derives from its vast number of learnable parameters, training on expansive corpora, and strong alignment with human preferences. Nevertheless, despite their potential to comprehend and design biomolecules in natural language following human intentions, leveraging LLMs for biomolecular research is still in its nascent stage and presents several challenges. Firstly, biomolecular research demands substantial specialized knowledge. General-purpose LLMs may not possess deep insights into this field because they are not tailored for specialized domains. Consequently, the lexical and semantic gap between the natural language and the language used to describe biomolecules presents its challenges. For instance, “C” could represent an alphabet letter in English, a carbon atom in a chemical molecule, or cysteine in the context of amino acids in proteins. Such semantic discrepancies might confuse models that are not specifically designed for these contexts. Secondly, unlike the primarily textual focus of general LLMs, biomolecules are intrinsically multimodal. For example, molecules are often represented by sequences like SMILES [9] or SELFIES [10]. Also, molecules can be inherently depicted as 2D graphs, featuring atoms as nodes and chemical bonds as edges, or as 3D structures, noting the spatial arrangements of atoms. Similarly, proteins are described through FASTA [11] sequences for their amino acid composition, and they also have 3D structures, essential for understanding their interactions and functions in space [12]. The effective utilization of this multimodal data presents a unique challenge for text-focused LLMs. Lastly, and most importantly, general LLMs struggle to align human intention in biomolecular tasks. To achieve the desired performance in a specific domain, LLMs need to be trained with alignment to acquire task-specific knowledge and patterns [13]. Instruction plays a critical role in this alignment process. By providing carefully curated, domain-specific instructions, LLMs are guided to develop a deeper understanding and more precise execution of specific tasks. In particular, to follow instructions for biomolecules, mastering the alignment between natural language, molecules, and proteins is essential. Complex tasks such as designing molecules for target proteins or creating enzymes for substrates necessitate simultaneous processing of natural language, molecules, and proteins.
@@
Although several recent endeavors [14, 15, 16, 17, 18, 19] have sought to tailor the LLMs for biomolecular tasks via extensive instructions, they encounter two primary limitations: (1) They typically align natural language with either molecules or proteins, but not both, lacking any-to-any alignment. (2) They also fall short in processing multimodal biomolecules, failing to align the multimodal data with natural language.
@@
Figure 1: An overview of InstructBioMol. a, InstructBioMol is a unified multimodal Large Language Model for natural language, molecules, and proteins. It can accept inputs in the form of natural language text, multimodal molecule and protein data and generate outputs as natural language, molecules, or proteins in the textual form. b, The Motif-Guided Multimodal Feature Extraction Module processes 2D graphs and 3D structures of molecules, as well as 1D sequences and 3D structures of proteins. Pre-trained modality-specific encoders obtain representations from these inputs, which are then processed by a Transformer Encoder. The Transformer Decoder, using motif prompts and learnable queries, produces multimodal features for integration into the language model. c, We collect datasets on a hundred-million scale, categorized into continual pretraining data and instruction-tuning data. Instruction-tuning is used to achieve an any-to-any alignment between molecule, protein, and natural language. A two-stage instruction-tuning paradigm enables the model to learn from low-quality extensive data (stage-1) to high-quality refined data (stage-2). d, InstructBioMol achieves alignment between molecules and natural language, proteins and natural language, as well as molecules and proteins. This enables it to follow human instructions, facilitating the understanding and design of biomolecules.
@@ Results
InstructBioMol can understand and design molecules following human intention
@@
We evaluate InstructBioMol’s capability in understanding and designing molecules through two tasks: (1) molecule captioning, which involves generating a textual description for a molecule; and (2) description-based molecule generation, where a molecule is generated based on a provided textual description. These tasks are introduced by ref. 14, and use the ChEBI dataset [22], which contains molecules and their corresponding descriptions including structure, function, origin, etc. Following ref. 14, for the first task, we select BLEU [23], ROUGE [24], and METEOR [25] as evaluation metrics. For the second task, we employ EXACT to measure the exact match of generated molecules, and using a range of similarity metrics — including BLEU, LEVENSHTEIN [26], MACCS FTS [27], RDK FTS [28], MORGAN FTS [29], and FCD [30] — to assess the similarity between generated molecules and ground truth. Furthermore, we utilize VALIDITY to evaluate the chemical validity of the generated molecules. Details of evaluation metrics are described in Methods. In this experiment, we evaluate two types of baselines. (1) Generalist language models. We assess two variants of general-purpose models. First, GPT-3.5 (zero-shot), a commonly used Large Language Model. Second, GPT-3.5 (10-shot MolReGPT) and GPT-4 (10-shot MolReGPT), which are adaptations of GPT-3.5 and GPT-4, respectively, using MolReGPT’s few-shot in-context learning approach [31]. (2) Molecule-specific enhanced language models. These models are further fine-tuned for molecular tasks, building on generalist models. Baselines include ChemDFM [32], InstructMol [33], MolT5 [14], BioT5 [16], and BioT5+ [18]. Notably, InstructMol is incapable of molecule generation.
@@ Results.
Quantitative results on molecule captioning and molecule generation are in Table 1 and Table 2, respectively. The performance of molecule-specific models significantly surpasses that of generalist language models, primarily due to the finetuning of the latter on domain-specific instruction datasets. This finetuning facilitates an effective alignment between natural language and chemical molecular knowledge. These findings suggest that general language models lack sufficient expertise in specialized domains, which can be effectively compensated by leveraging instruction-tuning. Notably, the experimental results demonstrate that InstructBioMol performs best across almost all evaluation metrics. Specifically, for the molecule captioning task, InstructBioMol yields an average improvement of 0.9% across all metrics. In description-based molecule generation task, the exact match accuracy (EXACT) of generated molecules increases by 0.7%. Furthermore, an average improvement of 2.0% is observed in molecular fingerprint similarity metrics (MACCS FTS, RDK FTS, and MORGAN FTS). These results indicate that InstructBioMol exhibits higher accuracy and efficacy in both understanding and generating chemical molecular information. We attribute this success to the extensive use of high-quality instruction data, which enables the model to comprehensively align molecules and natural language and achieve superior performance across molecular tasks. Some examples in Supplementary Information Section 4.1 provide a detailed analysis of the results generated by InstructBioMol.
Table 1: Performance comparison on molecule captioning task. (↑) / (↓) denotes a higher / lower value is better. The best performance is marked as bold.
@@
Table 2: Performance comparison on description-based molecule generation task. (↑) / (↓) denotes a higher / lower value is better. The best performance is marked as bold.
@@
Figure 2: Model performance on protein understanding and design benchmarks. a, Regarding protein understanding, models are evaluated by examining their performance in answering questions related to protein family, subcellular location, name, and function. The evaluation metrics employed include BLEU-2, BLEU-4, ROUGE-1, ROUGE-2, ROUGE-L, and METEOR. b, For description-based protein generation task, the accuracy and biological validity of the generated proteins are assessed using metrics including Identity, Alignment, BLOSUM Substitution, and Validity.
InstructBioMol can understand and design proteins following human intention
@@
We evaluate the model’s ability to understand and design proteins in the following two tasks: (1) answering questions about the properties of proteins, including protein family, subcellular location, official name, and function; (2) generating protein sequences based on the textual descriptions. For the protein property answering task, we apply the same evaluation metrics as for the molecule captioning task to assess the similarity between the generated answers and ground truth. For the description-based protein generation task, we use Identity, Alignment, and BLOSUM Substitution to measure the similarity of the generated proteins to ground truth. In particular, Identity captures exact amino acid matches; Alignment assesses the similarity of subsequences between proteins; and BLOSUM Substitution evaluates potential evolutionary relevance through amino acid substitutions. Higher scores on these three metrics indicate that the generated sequence is more similar to the ground truth. We also report the Validity for evaluating the biological validity of protein generation. These metrics are described in Methods in detail. In the experiments, we compare the following baseline methods: (1) the general-purpose language model GPT-3.5. GPT-3.5 is evaluated in three variants: GPT-3.5 (zero-shot), GPT-3.5 (5-shot, random), and GPT-3.5 (5-shot, similarity). The GPT-3.5 (5-shot, random) and GPT-3.5 (5-shot, similarity) variants utilize an in-context learning paradigm (detailed in Supplementary Information), where the former randomly selects 5 examples from the training set as prompts, and the latter selects the 5 most similar examples to the given query. (2) Protein-specific enhanced language models, which extend general language models to protein-related tasks, including Mol-Instructions [17], InstructProtein [15], BioT5+ [18], ProtT3 [34], BioMedGPT [19], and ProteinDT [35]. However, among these models, BioMedGPT and ProtT3 cannot generate protein sequences, while ProteinDT cannot answer protein property-related questions using natural language.
@@
Quantitive results on answering protein properties and description-based protein generation are in Figure 2a and Figure 2b, respectively. Based on the experimental results, we have reached the following conclusions: Firstly, InstructBioMol demonstrates the best performance in both tasks. In tasks related to answering questions about protein properties, InstructBioMol outperforms previous state-of-the-art (SOTA) methods by 13.1% on average. For protein generation tasks, InstructBioMol achieves 0.9%, 5.7%, and 0.510 improvements in identity, alignment, and BLOSUM substitution, respectively, compared to previous SOTAs, and comparable validity of the generated proteins. Secondly, using domain-specific instruction alignment is effective. In tasks related to answering protein property questions, InstructBioMol shows an average improvement of 80.3% compared to GPT-3.5 (zero-shot). In protein generation tasks, InstructBioMol achieved a 17.5% increase in identity and a 28.7% increase in alignment compared to GPT-3.5 (zero-shot). These results clearly demonstrate that models aligned with domain-specific instructions exhibit significantly enhanced capabilities in handling tasks within specific domains compared to general models. This underscores the importance of customized models in specialized domains, particularly in highly specialized fields like protein engineering, where the integration of domain-specific knowledge and instructions can greatly enhance the model’s practicality and accuracy. In Supplementary Information Section 4.2, we provide a detailed analysis of InstructBioMol’s outstanding performance on protein function answering tasks through some examples. Additionally, example cases on description-based protein generation task demonstrate that InstructBioMol can design de-novo proteins with high structural similarity, closely aligning with the structures of ground truth.
Figure 3: Model performance on drug discovery and enzyme design. a-c, Performance comparison on drug discovery. (a), The generated drug-like molecules are evaluated from multiple perspectives, including binding affinity (e.g., Vina Score and High Affinity), general properties (SA, QED, and Validity), and an overall evaluation metric (Success Rate). b-c, A detailed analysis of Vina Scores is presented, including (b) top-1, top-5, top-10, and all Vina Scores, as well as (c) the distribution of Vina Scores for all generated molecules. d, Performance comparison on enzyme design. Evaluation metrics include similarity metrics Identity and Alignment; interaction metrics ESP Score and Vina Score; and Validity.
InstructBioMol enables target protein-based drug discovery following human intention
@@
Designing molecules that can bind to specific proteins is one of the most challenging tasks in drug discovery [36]. The chemical space is vast, yet the subset of molecules with desirable biological activity is relatively small. Drug discovery typically involves searching this expansive space for molecules that can bind to specific targets, such as disease-related proteins. As a research copilot, InstructBioMol can design molecule drugs for target proteins following human intention, thereby reducing the time and cost of drug development. For experimental validation, we have established a series of evaluation metrics. Following refs.37, 38, evaluation metrics can be divided into three perspectives, including (1) quantitative estimates of binding affinity: Vina Score [39] and High Affinity. Vina Score reflects the best possible binding affinity, where a lower score indicating stronger affinity. High Affinity measures the ratio of generated molecules that bind stronger than the reference molecules in the test set. (2) General molecular properties: synthetics accessibility (SA) [40], drug-likeness (QED) [41], and Validity. (3) Overall assessment following ref. 38: Success Rate, which measures the ratio of molecules with binding affinity versus favorable properties (Vina Score $<$ -8.18, QED $>$ 0.25, SA $>$ 0.59). Details of metrics are described in Methods. To make a comparison, we take three variants of GPT-3.5 (zero-shot, 5-shot random, and 5-shot similarity) and DrugGPT [42] as baselines, and also take the molecules in the test set as the Reference baseline. 100 target proteins are selected in the test set, and for each target protein, 100 molecules are generated.
@@
Figure 3a presents the experimental results. where InstructBioMol demonstrates superior performance across three key dimensions: binding affinity, general properties, and overall assessment. Specifically, it improves High Affinity and Success Rate by 25.9% and 21.9%, respectively, compared to previous state-of-the-art (SOTA) methods. This highlights InstructBioMol’s enhanced capability in generating drug-like molecules with high affinity for target proteins and favorable intrinsic properties. Additionally, it achieves an outstanding generation validity of 99.9%. Figure 3b illustrates the average Vina Scores for top-1, top-5, top-10, and all generated molecules. It can be observed that InstructBioMol consistently outperforms other methods under these settings, and is the only approach where the average scores for all generated molecules surpass the reference values. This suggests that the quality of the molecules generated by InstructBioMol is comparable to the ground truth in the dataset. Moreover, InstructBioMol proves to be the most effective method for designing molecules with the best Vina Scores for most target proteins. (As shown in Figure 4a in Supplementary Information, InstructBioMol achieves the best performance on 35% of the targets.) Figure 3c shows the distribution of Vina Scores for all generated molecules. The distribution reveals that molecules generated by InstructBioMol have a lower mean and reduced variance, further confirming that the overall quality of these molecules is superior to that of other methods.
Figure 4: Results of ablation analysis on the impact of training strategy and the multimodal data integration. a, The impact of removing specific training stages is evaluated across four tasks: molecule captioning, description-based molecule generation, protein function answering, and description-based protein generation. The models compared include InstructBioMol and its two variants: one variant retains the continual-pretraining stage but removes the stage-1 instruction-tuning (w/ continual-pretraining, w/o stage-1 instruction-tuning), while the other variant removes both the continual-pretraining stage and the stage-1 instruction-tuning (w/o continual-pretraining, w/o stage-1 instruction-tuning). b, The comparison of multimodal data integration, involving InstructBioMol and its various variants: w/o 1D encoder, w/o 2D encoder, and w/o 3D encoder, which correspond to the removal of specific modality inputs. Additionally, w/o motif denotes the removal of motif prompts, and w/o multimodal represents the removal of the entire Motif-Guided Multimodal Feature Extraction Module.
InstructBioMol enables target substrate-based enzyme design following human intention
@@
Enzymes, as biological catalysts, can accelerate chemical reactions in various biological processes [43]. In enzymatic reactions, substrates are the small molecules that are catalytically converted by enzymes. By binding to specific substrates and acting upon them, enzymes significantly enhance the conversion efficiency of the substrates. Designing enzymes that can bind to specific substrates is a crucial and challenging research problem. InstructBioMol can assist researchers in designing protein enzymes for specific substrates, thereby advancing the progress of efficient enzyme design. To establish experimental validation, we split a test set containing 100 substrates, and for each target substrate, 100 enzyme proteins are generated. The following evaluation metrics are employed, which are divided into three groups: (1) similarity between generated proteins and ground truth, including Identity and Alignment. (2) Substrate-enzyme interaction, assessed using ESP Score [20] and Vina Score. ESP Score measures the interaction capability of the designed enzymes to their substrates, with a range of 0-100, where higher scores indicate stronger interactions. Vina Score quantifies the strength of binding affinity between the designed enzymes and substrates, with lower scores indicating higher affinity. (3) Validity is used to assess whether generated enzymes are biologically valid. Details of metrics are in Methods. For comparison, we employ three variants of GPT-3.5 (zero-shot, 5-shot random, 5-shot similarity) and RFdiffusionAA [5] as baseline models.

### v2

# Advancing Biomolecular Understanding and Design Following Human Instructions

@@ Abstract
Understanding and designing biomolecules, such as proteins and small molecules, is central to advancing drug discovery, synthetic biology, and enzyme engineering. Recent breakthroughs in Artificial Intelligence (AI) have revolutionized biomolecular research, achieving remarkable accuracy in biomolecular prediction and design. However, a critical gap remains between AI’s computational capabilities and researchers’ intuitive goals, particularly in using natural language to bridge complex tasks with human intentions. Large Language Models (LLMs) have shown potential to interpret human intentions, yet their application to biomolecular research remains nascent due to challenges including specialized knowledge requirements, multimodal data integration, and semantic alignment between natural language and biomolecules. To address these limitations, we present InstructBioMol, a novel LLM designed to bridge natural language and biomolecules through a comprehensive any-to-any alignment of natural language, molecules, and proteins. This model can integrate multimodal biomolecules as input, and enable researchers to articulate design goals in natural language, providing biomolecular outputs that meet precise biological needs. Experimental results demonstrate InstructBioMol can understand and design biomolecules following human instructions. Notably, it can generate drug molecules with a 10% improvement in binding affinity and design enzymes that achieve an Enzyme-Substrate Pair Prediction Score (ESP Score) of 70.4, making it the only method to surpass the enzyme-substrate interaction threshold of 60.0 recommended by the ESP developer. This highlights its potential to transform real-world biomolecular research.
@@ Introduction
Understanding and designing biomolecules is fundamental to natural science research. Biomolecules, such as proteins and small molecules, play essential roles in biological processes, and their precise manipulation is key to advancements in drug discovery, synthetic biology, and enzyme engineering [1, 2, 3]. Recent Artificial Intelligence (AI) breakthroughs have transformed research in these areas [4, 5]. Tools like AlphaFold3 [4] and RoseTTAFold All-Atom [5] have revolutionized biomolecular structure prediction, offering unprecedented accuracy and speed. Despite these advancements, a crucial challenge persists: how to effectively understand biomolecules using natural language and design them according to human intentions. This presents a gap between AI’s computational capacity and researchers’ needs to apply it to real-world problems. Consider the scenario of a researcher tasked with designing a new drug to target a protein involved in a complex disease. Traditionally, this process involves navigating vast amounts of biomolecular data, interpreting biological and chemical relationships, and iterating through trial-and-error to engineer molecules with specific properties. While AI has enhanced many aspects of this workflow, current tools often struggle to align molecular complexity with intuitive, human-driven goals articulated in natural language.
@@
Notably, Large Language Models (LLMs) [6] have demonstrated their impressive capacity to understand human intention and generate human-like responses helpfully and safely [7, 8]. This capability derives from its vast number of learnable parameters, training on expansive corpora, and strong alignment with human preferences. Nevertheless, despite their potential to comprehend and design biomolecules in natural language following human intentions, leveraging LLMs for biomolecular research is still in its nascent stage and presents several challenges [9]. Firstly, biomolecular research demands substantial specialized knowledge. General-purpose LLMs may not possess deep insights into this field because they are not tailored for specialized domains. Consequently, the lexical and semantic gap between the natural language and the language used to describe biomolecules presents its challenges. For instance, “C” could represent an alphabet letter in English, a carbon atom in a chemical molecule, or cysteine in the context of amino acids in proteins. Such semantic discrepancies might confuse models that are not specifically designed for these contexts. Secondly, unlike the primarily textual focus of general LLMs, biomolecules are intrinsically multimodal. For example, molecules are often represented by sequences like SMILES [10] or SELFIES [11]. Also, molecules can be inherently depicted as 2D graphs, featuring atoms as nodes and chemical bonds as edges, or as 3D structures, noting the spatial arrangements of atoms. Similarly, proteins are described through FASTA [12] sequences for their amino acid composition, and they also have 3D structures, essential for understanding their interactions and functions in space [13]. The effective utilization of this multimodal data presents a unique challenge for text-focused LLMs. Lastly, and most importantly, general LLMs struggle to align human intention in biomolecular tasks. To achieve the desired performance in a specific domain, LLMs need to be trained with alignment to acquire task-specific knowledge and patterns [14]. Instruction plays a critical role in this alignment process. By providing carefully curated, domain-specific instructions, LLMs are guided to develop a deeper understanding and more precise execution of specific tasks. In particular, to follow instructions for biomolecules, mastering the alignment between natural language, molecules, and proteins is essential. Complex tasks such as designing molecules for target proteins or creating enzymes for substrates necessitate simultaneous processing of natural language, molecules, and proteins.
@@
Although several recent endeavors [15, 16, 17, 18, 19, 20, 21] have sought to tailor the LLMs for biomolecular tasks via extensive instructions, they encounter two primary limitations: (1) They typically align natural language with either molecules or proteins, but not both, lacking any-to-any alignment. (2) They also fall short in processing multimodal biomolecules, failing to align the multimodal data with natural language.
@@
Figure 1: An overview of InstructBioMol. a, InstructBioMol is a unified multimodal Large Language Model for natural language, molecules, and proteins. It can accept inputs in the form of natural language text, multimodal molecule and protein data and generate outputs as natural language, molecules, or proteins in the textual form. b, The Motif-Guided Multimodal Feature Extraction Module processes 2D graphs and 3D structures of molecules, as well as 1D sequences and 3D structures of proteins. Pre-trained modality-specific encoders obtain representations from these inputs, which are then processed by a Transformer Encoder. The Transformer Decoder, using motif prompts and learnable queries, produces multimodal features for integration into the language model. The 2D molecule illustration is derived using SmilesDrawer [22], and the 3D molecule and protein illustrations are from PyMOL [23]. c, We collect datasets on a hundred-million scale, categorized into continual pretraining data and instruction-tuning data. Instruction-tuning is used to achieve an any-to-any alignment between molecule, protein, and natural language. A two-stage instruction-tuning paradigm enables the model to learn from low-quality extensive data (stage-1) to high-quality refined data (stage-2). d, InstructBioMol achieves alignment between molecules and natural language, proteins and natural language, as well as molecules and proteins. This enables it to follow human instructions, facilitating the understanding and design of biomolecules. The icons are from Flaticon.com.
@@ Results
InstructBioMol can understand and design molecules
@@
We evaluate InstructBioMol’s capability in understanding and designing molecules through two tasks: (1) molecule captioning, which involves generating a textual description for a molecule; and (2) description-based molecule generation, where a molecule is generated based on a provided textual description. These tasks are introduced by ref. 15, and use the ChEBI dataset [26], which contains molecules and their corresponding descriptions including structure, function, origin, etc. The evaluation metrics are consistent with those in ref. 15 and described in Methods. In this experiment, we evaluate two types of baselines. (1) Generalist language models. We assess two variants of general-purpose models. First, GPT-3.5 (zero-shot), a commonly used Large Language Model. Second, GPT-3.5 (10-shot MolReGPT) and GPT-4 (10-shot MolReGPT), which are adaptations of GPT-3.5 and GPT-4, respectively, using MolReGPT’s few-shot in-context learning approach [27]. (2) Molecule-specific enhanced language models. These models are further fine-tuned for molecular tasks, building on generalist models. Baselines include ChemDFM [28], InstructMol [29], MolT5 [15], BioT5 [17], and BioT5+ [19].
@@ Results.
Quantitative results on molecule captioning and molecule generation are in Table 1 and Table 2, respectively. Molecule-specific models outperform generalist language models due to finetuning on domain-specific instruction datasets, which align natural language with chemical knowledge. This highlights that general models lack domain expertise, which can be addressed through instruction-tuning. Notably, the experimental results demonstrate that InstructBioMol performs best across almost all evaluation metrics. Specifically, for the molecule captioning task, InstructBioMol yields an average improvement of 0.9% across all metrics. In description-based molecule generation task, the exact match accuracy (EXACT) of generated molecules increases by 0.7%. Furthermore, an average improvement of 2.0% is observed in molecular fingerprint similarity metrics (MACCS FTS, RDK FTS, and MORGAN FTS). These results indicate that InstructBioMol exhibits higher accuracy and efficacy in both understanding and generating chemical molecular information. We attribute this success to the extensive use of high-quality instruction data, which enables the model to comprehensively align molecules and natural language and achieve superior performance across molecular tasks.
Table 1: Performance comparison on molecule captioning task. (↑) / (↓) denotes that a higher / lower value is better. The best performance is marked in bold.
@@
Table 2: Performance comparison on description-based molecule generation task. (↑) / (↓) denotes that a higher / lower value is better. The best performance is marked in bold.
@@
Figure 2: Model performance on protein understanding and design benchmarks. a, Regarding protein understanding, models are evaluated by examining their performance in answering questions related to protein family, subcellular location, name, and function. The evaluation metrics employed include BLEU-2, BLEU-4, ROUGE-1, ROUGE-2, ROUGE-L, and METEOR. b, For description-based protein generation task, the accuracy and biological validity of the generated proteins are assessed using metrics including Identity, Alignment, BLOSUM Substitution, and Validity. The number of experimental replicates is $n$=3, and the error bars are standard deviation.
Figure 3: Model performance on drug discovery and enzyme design. a-c, Performance comparison on drug discovery. (a), The generated drug-like molecules are evaluated from multiple perspectives, including binding affinity (e.g., Vina Score and High Affinity), general properties (SA, QED, and Validity), and an overall evaluation metric (Success Rate). b-c, A detailed analysis of Vina Scores is presented, including (b) top-1, top-5, top-10, and all Vina Scores, as well as (c) the distribution of Vina Scores for all generated molecules. ($n$=10,000, median as centre line, 25th to 75th percentiles as bounds of the box, whiskers extending to 1.5 times the interquartile range from the bounds of the box.) d, Performance comparison with statistical significance on enzyme design. Evaluation metrics include similarity metrics Identity and Alignment; interaction metrics ESP Score and Vina Score; and Validity. Statistical significance is determined by two-sided independent sample t-test.
InstructBioMol can understand and design proteins
@@
We evaluate the model’s ability to understand and design proteins in the following two tasks: (1) answering questions about the properties of proteins, including protein family, subcellular location, official name, and function; (2) generating protein sequences based on the textual descriptions. The evaluation metrics are described in Methods. In the experiments, we compare the following baseline methods: (1) the general-purpose language model GPT-3.5. GPT-3.5 is evaluated in three variants: GPT-3.5 (zero-shot), GPT-3.5 (5-shot, random), and GPT-3.5 (5-shot, similarity). The GPT-3.5 (5-shot, random) and GPT-3.5 (5-shot, similarity) variants utilize an in-context learning paradigm (detailed in Supplementary Information Section 2), where the former randomly selects 5 examples from the training set as prompts, and the latter selects the 5 most similar examples to the given query. (2) Protein-specific enhanced language models, which extend general language models to protein-related tasks, including Mol-Instructions [18], InstructProtein [16], BioT5+ [19], ProtT3 [30], BioMedGPT [20], ProteinDT [31], and ProteinDT+motif, a variant of ProteinDT that incorporates motif embeddings during training.
@@
Quantitative results on answering protein properties and description-based protein generation are in Figure 2a and Figure 2b, respectively. Based on the experimental results, we have reached the following conclusions: Firstly, InstructBioMol demonstrates the best performance in both tasks. In tasks related to answering questions about protein properties, InstructBioMol outperforms previous state-of-the-art (SOTA) methods by 13.5% on average. For protein generation tasks, InstructBioMol achieves 6.2%, 4.1%, and 0.465 improvements in Identity, Alignment, and BLOSUM Substitution, respectively, compared to previous SOTAs, and comparable validity of the generated proteins. Secondly, domain-specific instruction alignment proves highly effective. In protein property question tasks, InstructBioMol outperforms GPT-3.5 (zero-shot) significantly, and in protein generation, it shows a 17.8% increase in Identity and a 27.7% improvement in Alignment. These results highlight that models tailored with domain-specific instructions significantly outperform general models in specialized tasks, particularly in fields like protein engineering, where integrating domain knowledge greatly enhances both practicality and accuracy.
InstructBioMol enables target protein-based drug discovery
@@
Designing molecules that can bind to specific proteins is one of the most challenging tasks in drug discovery [32]. The chemical space is vast, yet the subset of molecules with desirable biological activity is relatively small. Drug discovery typically involves searching this expansive space for molecules that can bind to specific targets, such as disease-related proteins. As a research copilot, InstructBioMol can design molecule drugs for target proteins following human intention, thereby reducing the time and cost of drug development. In this experiment, evaluation metrics are detailed in Methods. To make a comparison, we take three variants of GPT-3.5 (zero-shot, 5-shot random, and 5-shot similarity), Pocket2Mol [33], AR-SBDD [34], TargetDiff [35] and DrugGPT [36] as baselines, and also take molecules in the test set as the Reference baseline. 100 target proteins are selected in the test set. For each target protein, 100 molecules are generated.
@@
Figure 3a presents the experimental results. where InstructBioMol demonstrates superior performance across three key dimensions: binding affinity, general properties, and overall assessment. Specifically, it improves High Affinity and Success Rate by 25.9% and 21.9%, respectively, compared to previous state-of-the-art (SOTA) methods. This highlights InstructBioMol’s enhanced capability in generating drug-like molecules with high affinity for target proteins and favorable intrinsic properties. Additionally, it achieves an outstanding generation validity of 99.9%. Figure 3b illustrates the average Vina Scores for top-1, top-5, top-10, and all generated molecules. It is evident that InstructBioMol consistently outperforms other methods in these settings, achieving a $P$-value of 0.05 for outperforming the second-best model (DrugGPT) in the top-1 Vina Score. Moreover, it is the only approach where the average scores for all generated molecules exceed the reference values. This suggests that the quality of the molecules generated by InstructBioMol is comparable to the ground truth in the dataset. Moreover, InstructBioMol proves to be the most effective method for designing molecules with the best Vina Scores for most target proteins. As shown in Supplementary Figure 1a, InstructBioMol achieves the best performance on 35% of the targets. Figure 3c shows the distribution of Vina Scores for all generated molecules. The distribution reveals that molecules generated by InstructBioMol have a lower mean and reduced variance, further confirming that the overall quality of these molecules is superior to that of other methods.
InstructBioMol enables target substrate-based enzyme design
@@
Enzymes, as biological catalysts, can accelerate chemical reactions in various biological processes [37]. In enzymatic reactions, substrates are the small molecules that are catalytically converted by enzymes. By binding to specific substrates and acting upon them, enzymes significantly enhance the conversion efficiency of the substrates. Designing enzymes that can bind to specific substrates is a crucial and challenging research problem. InstructBioMol can assist researchers in designing protein enzymes for specific substrates, thereby advancing the progress of efficient enzyme design. To establish experimental validation, we split a test set containing 100 substrates, and for each target substrate, 100 enzyme proteins are generated. The evaluation metrics employed in this experiment are detailed in Methods. For comparison, we employ three variants of GPT-3.5 (zero-shot, 5-shot random, 5-shot similarity) and RFdiffusionAA [5] as baselines.